# Ultracold Interactions between Ions and Polar Molecules

### v1

@@ 3 Polar molecules in the electric field of an atomic ion
In addition to the difficulty of reaching the required extremely low collision energies, another effect potentially poses an even stricter limitation on the range of observable effects: the enhancement of the loss processes due to three-body recombination. From ion-atom collisions it is known that ion-atom-atom TBR becomes the dominant loss process already at comparatively low densities several orders of magnitude below values where atomic ensembles suffer from three-body losses [23], even in the case of spin-polarized fermions [17]. This can be understood intuitively by coarsely estimating the critical density $n_{C}$ at which the interparticle volume becomes smaller than the active volume $V^{\star}=(4/3)\pi R^{\star}$3. The latter is determined by $R^{\star}(n)=\left(n\mu C_{n}/(2\hbar)\right)^{1/(n-2)}$ of the dominant interaction $-C_{n}/R^{n}$ [79]. For $n>n_{C}=V^{\star}$-1 more than one particle enter $V^{\star}$ at the same time, marking the onset of TBR. In the case of ground-state neutral atoms, $R^{\star}$ arises from the short-range van der Waals potential $-C_{6}/R^{6}$. For example, in rubidium $R^{\star}\approx 10~{}\textrm{nm}$ and $n_{C}=2.5\times 10^{17}~{}\textrm{cm}^{-3}$, about two orders of magnitude higher than typical BEC densities. For ground-state atoms colliding with ions, e.g. Rb and Ba+, $R^{\star}\approx 290~{}\textrm{nm}$ of the charge-induced dipole interaction $-C_{4}/R^{4}$ is about 30 times larger. Therefore, the critical density is already reached at $n_{C}^{Ba-Rb}\approx 1\times 10^{13}~{}\textrm{cm}^{-3}$ which is in agreement with experimental findings [23, 25, 24, 21]. For Ba ions interacting with NaK molecules, $R^{\star}$ is once again drastically increased to $R^{\star}=1.42\times 10^{-5}~{}\textrm{m}$, such that $V^{\star}$ would be comparable to the capture volume of typical dipole traps and $n_{C}^{Ba-NaK}\approx 1\times 10^{8}~{}\textrm{cm}^{-3}$.
Following the analogy to ion-atom interactions, it becomes apparent that reaching the quantum dominated regime in the Ba+-NaK system using even the most advanced hybrid platforms is extremely challenging. However, as is the case with ion-atom systems, these results strongly depend on the reduced masses and static polarizabilities of the chosen ion-molecule combination. For example, for Ba+-LiNa, the situation is much more comparable to Ba+-Rb, with both $E^{\star}$ and $R^{\star}$ being of the same order of magnitude.
In addition to the choice of a suitable combination, all the characteristic properties that are fixed for a given combination of ions and ground-state atoms can be readily and drastically tuned in polar molecules, including those with an extremely low threshold for s-wave collisions such as NaK and NaCs. Since the polarizability in the relevant limit is simply determined by the second derivative of the potential with respect to $\mathcal{E}$, we are free to choose any addressable rotational state which will exhibit not only a different magnitude of the polarizability but for particular projections $m_{J}$ also an opposite character of the polarization potential. For instance, already the first rotationally excited state $\left|X^{1}\Sigma^{+},v=0,J=1,m_{J}=0\right\rangle$ of NaK features a reversed curvature as shown in Fig.4(b). This means that the polarization potential is rendered repulsive and features a barrier with a height of $U_{0}^{(J=1,m_{J}=0)}/k_{B}=87.4~{} \textrm{mK}$ reached around $\mathcal{E}=10~{}\textrm{kV cm}^{-1}\approx 5~{}\mathcal{E}^{\star}$ in NaK as depicted in Fig.3. Therefore, for molecules prepared in this state and at typically achievable temperatures on the order of $T_{mol}\sim 100~{}\textrm{nK}\ll U_{0}^{(J=1,m_{J}=0)}/k_{B}$, we expect that short-range ion-molecule collisions will be strongly suppressed.
@@
In ion-atom systems, conceptually similar ideas for designing effectively repulsive potentials by exploiting Rydberg dressing induced by optical fields have been proposed [31]. In this case, a potential barrier arises from the competition of an engineered repulsive contribution and the attractive static charge-induced dipole potential, both scaling as $R^{-4}$. In contrast, for molecules interacting with ions, the long range potential itself is repulsive such that the sought-after shielding can be facilitated by exploiting their rich energy level structure, that is, without additional dressing.
This precise control over the rotational degrees of freedom provides a natural mechanism for shielding against inelastic and reactive ion-molecule collisions, opening a way to realize sympathetic cooling. We note, that this expectation holds even in presence of ion micromotion and despite the prediction that $E^{\star}$ in the rovibronic ground state is extremely low compared to the thermal energy of typical molecular quantum ensembles. This intriguing feature hints at the fact that the intuition derived from considering characteristic quantities such as $R^{\ast}$ and $E^{\ast}$ is only valid for the specific case of molecules prepared in the rovibronic ground state. Furthermore, even in the latter case, the comparison to ion-atom systems only holds in the regime of extremely dilute gases where molecules collide with the ion one at a time, such that the intermolecular interactions are negligible. In contrast, this approximation breaks down for dense molecular gases where the dipole-dipole interactions and many-body effects are expected to play a significant role [80]. This scenario is discussed in section 5.
4 Reactive collisions and formation of molecular ions

### v2

@@ II.1 Molecules in the rovibronic ground state
Table 1: Characteristic properties of the charge-induced dipole long-range interaction between common heteronuclear bialkali and laser-coolable molecules prepared in the rovibronic ground state and a 138Ba+ ion. The static dipole polarizabilities $\alpha_{0}$ and permanent electric dipole moments (PDM) of the bialkali molecules are taken from [60]. The polarizabilities due to purely rotational transitions of the laser-coolable species were approximated as $\alpha_{0}^{rot}\approx d_{0}^{2}/(3B_{0})$ [64] with constants $d_{0}$ and $B_{0}$ for CaH [65, 66], CaF [67], CaOH [68, 69]. For comparison, 87Rb and 6Li are given as a references for Ba+-atom systems, representing a typical and one of the most favorable combinations with respect to reaching the s-wave scattering limit, respectively.
PDM $B_{0}/h$ $\alpha_{0}$ $R^{\star}$ $E^{\star}/k_{\textrm{B}}$
(D) (GHz) $(10^{-37}~{}\textrm{C m}^{2}\textrm{V}^{-1})$ $(10^{-6} \textrm{m})$ (nK)
LiCs $5.59$ $5.61$ $312$ $25.9$ $0.00521$
NaCs $4.69$ $1.74$ $709$ $40.1$ $0.00206$
LiRb $4.18$ $6.45$ $152$ $16.2$ $0.0165$
LiK $3.58$ $7.67$ $93.6$ $10.0$ $0.0703$
NaRb $3.31$ $2.10$ $294$ $23.6$ $0.00711$
NaK $2.78$ $2.85$ $152$ $14.2$ $0.0281$
KCs $1.84$ $0.899$ $208$ $22.2$ $0.00644$
RbCs $1.25$ $0.48$ $177$ $21.6$ $0.00614$
KRb $0.62$ $1.14$ $18.8$ $6.21$ $0.0952$
LiNa $0.57$ $11.2$ $1.64$ $1.12$ $7.86$
CaH $2.94$ $128$ $3.77$ $1.92$ $2.08$
CaF $3.07$ $10.3$ $51.4$ $8.11$ $0.0893$
CaOH $1.47$ $10.0$ $51.4$ $8.01$ $0.0937$
Rb $0.00$ $0.00$ $0.0526$ $0.295$ $52.3$
Li $0.00$ $0.00$ $0.0270$ $0.0694$ $8760$
@@
In addition to the seeming difficulty of reaching the required extremely low collision energies, another effect potentially poses an even stricter limitation on the range of observable effects: the enhancement of the loss processes due to three-body recombination. From ion-atom collisions it is known that ion-atom-atom TBR becomes the dominant loss process already at comparatively low densities several orders of magnitude below values where atomic ensembles suffer from three-body losses [25], even in the case of spin-polarized fermions [19]. This can be understood intuitively by coarsely estimating the critical density $n_{C}$ at which the interparticle volume becomes smaller than the active volume $V^{\star}=(4/3)\pi R^{\star}$3. The latter is determined by $R^{\star}(p)=\left(p\mu C_{m}/(2\hbar)\right)^{1/(p-2)}$ of the dominant interaction $-C_{p}/R^{p}$ [63]. For $n>n_{C}=V^{\star}$-1 more than one particle enter $V^{\star}$ at the same time. At this point TBR becomes the dominant loss process according to the simplified model, while under realistic conditions, TBR losses are significant already at lower densities. In the case of ground-state neutral atoms, $R^{\star}$ arises from the short-range van der Waals potential $-C_{6}/R^{6}$. For example, in rubidium $R^{\star}\approx 10~{}\textrm{nm}$ and $n_{C}=2.5\times 10^{17}~{}\textrm{cm}^{-3}$, about two orders of magnitude higher than typical BEC densities. For ground-state atoms colliding with ions, e.g. Rb and Ba+, $R^{\star}\approx 300~{}\textrm{nm}$ of the charge-induced dipole interaction $U_{CID}=-C_{4}/R^{4}$ is about 30 times larger. Therefore, the critical density is already reached at $n_{C}^{Ba-Rb}\approx 1\times 10^{13}~{}\textrm{cm}^{-3}$ which is in agreement with experimental findings [25, 27, 26, 23]. For Ba ions interacting with NaK molecules, $R^{\star}$ is once again drastically increased to $R^{\star}=1.42\times 10^{-5}~{}\textrm{m}$, such that $V^{\star}$ would be comparable to the capture volume of typical dipole traps and $n_{C}^{Ba-NaK}\approx 1\times 10^{8}~{}\textrm{cm}^{-3}$.
Following the analogy to ion-atom interactions, it would seem obvious that reaching the quantum dominated regime in the Ba+-NaK system using even the most advanced hybrid platforms would be extremely challenging. However, it is important to note that these results strongly depend on the reduced masses and static polarizabilities of the chosen ion-molecule combination. Moreover, the Langevin model only holds in the regime of extremely dilute gases, that is $n\ll n_{C}$, where molecules collide with the ion one at a time, such that the intermolecular interactions are negligible. In contrast, this approximation breaks down for dense molecular gases where the dipole-dipole interactions and many-body effects are expected to play a significant role [70]. This scenario is discussed in Section V. Most importantly, the characteristic quantities derived so far only apply to molecules in the rovibronic ground state $\left|J=0,m_{J}=0\right\rangle$. Therefore, as a next step, we consider the influence of other rotational states on the collisional properties of ion-molecule mixtures.
II.2 Molecules in rotationally excited states
The effective long-range polarizability $\alpha_{0,J}^{eff}$ of molecules prepared in the rovibronic state $\left|X^{1}\Sigma^{+},v=0,J,m_{J}\right\rangle$ contains two main contributions, $\alpha_{0,J}^{rot}$ arising from purely rotational transitions, and $\alpha_{0,J}^{el}$, the contribution due to vibronic transtions to excited electronic states, while the component attributed to rovibrational transitions within the electronic ground state is typically small [60]. Since $\alpha_{0,J}^{rot}$ is determined by the second derivative of the potential with respect to $\mathcal{E}$, as shown in Eqn. 3, both $\alpha_{0,J}^{eff}$ and $E^{\star}_{J}$ strongly depend on the choice of the rotational state. Therefore, by exploiting radiatively long-lived rotational states, all the characteristic properties that are fixed for a given combination of ions and ground-state atoms could be tuned by several orders of magnitude in polar molecules, including those with an extremely low threshold for s-wave collisions such as NaK and NaCs. The expected impact on the properties of ion-molecule mixtures is illustrated in Table 2 for two exemplary combinations, Ba+-NaK and Ba+-LiNa. In the following, we identify several states that may be advantageous for ion-molecule collisions in the ultracold regime, characterized by negative, negligible or low curvatures with respect to $\cal{E}$ at low field which we obtain by numerically calculating $\partial^{2}U_{CD}(\mathcal{E})/\partial\mathcal{E}^{2}$. In the latter instances, $\alpha_{0,J}^{el}$ is sizable and has to be taken into account. This is in contrast to the rotational ground state where typically $\alpha_{0,0}^{el}\ll\alpha_{0,0}^{rot}$.
Table 2: Effective static polarizabilities $\alpha_{0,J}^{eff}$ and characteristic energies $E^{\star}_{J}$ for collisions with Ba+ for selected rotational states $\left|J,m_{J}\right\rangle$. The low-field seeking state $\left|J=1,m_{J}=0\right\rangle$ yields imaginary $E^{\star}_{J}$.
rot. state $\alpha_{0,J}^{eff}$ $E^{\star}_{J}/k_{\textrm{B}}$
$\left|J,m_{J}\right\rangle$ $(10^{-37}~{}\textrm{C m}^{2}\textrm{V}^{-1})$ (nK)
NaK $\left|0,0\right\rangle$ $152$ $0.0282$
$\left|1,0\right\rangle$ $-91.1$
$\left|1,\pm 1\right\rangle$ $45.6$ $0.0939$
$\left|2,\pm 2\right\rangle$ $21.7$ $0.197$
$\left|3,\pm 2\right\rangle$ $0.0578$ $74.1$
LiNa $\left|0,0\right\rangle$ $1.62$ $7.86$
$\left|1,0\right\rangle$ $-0.935$
$\left|1,\pm 1\right\rangle$ $0.487$ $26.512$
$\left|2,\pm 2\right\rangle$ $0.232$ $55.676$
$\left|3,\pm 2\right\rangle$ $0.0385$ $335$
$\left|5,\pm 1\right\rangle$ $0.00109$ $11900$
@@ II.2 Molecules in rotationally excited states
One interesting example from this class is the first rotationally excited state $\left|X^{1}\Sigma^{+},v=0,J=1,m_{J}=0\right\rangle$ which features a reversed curvature as shown in Fig.2(b) for NaK. This means that the polarization potential is rendered repulsive with a barrier height of $U_{0}^{(J=1,m_{J}=0)}/k_{B}=87.4~{} \textrm{mK}$ reached around $\mathcal{E}=10~{}\textrm{kV cm}^{-1}\approx 5~{}\mathcal{E}^{\star}$ in NaK as depicted in Fig.1. Therefore, for molecules prepared in this state and at typically achievable temperatures on the order of $T_{mol}\sim 100~{}\textrm{nK}\ll U_{0}^{(J=1,m_{J}=0)}/k_{B}$, we expect that short-range ion-molecule collisions, and ion-molecule-molecule TBR, will be strongly suppressed, while many desired effects such as sympathetic cooling are still expected to be efficient. We note, that this expectation holds even in presence of ion micromotion and despite the prediction that $E^{\star}$ in the rovibronic ground state is extremely low compared to the thermal energy of typical molecular quantum ensembles. In ion-atom systems, conceptually similar ideas for designing effectively repulsive potentials by exploiting Rydberg dressing induced by optical fields have been proposed [33]. In this case, a potential barrier arises from the competition of an engineered repulsive contribution and the attractive static charge-induced dipole potential, both scaling as $R^{-4}$. In contrast, for molecules interacting with ions, the long-range potential itself is repulsive such that the sought-after shielding can be facilitated by exploiting their rich energy level structure, that is, without additional dressing.
As a second example we consider LiNa prepared in the state $\left|J=5,m_{J}=\pm 1\right\rangle$. We find that the polarizability contribution calculated from its curvature is opposite in sign and comparable in magnitude to that stemming from coupling to electronically excited states [60]. The effective polarizability is then about $10^{3}$ times smaller compared to $\left|J=0,m_{J}=0\right\rangle$, with the corresponding $E_{J}^{\star}/k_{B}\approx 10~{}\mu\text{K}$ being comparable to the most favorable ion-atom mixtures in rf-hybrid traps. For NaK, we find no states with such properties for rotational manifolds up to $J=7$. This suggests that entering the quantum regime of ion-molecule collisions may be feasible, with methodology similar to that established in ion-atom systems, for suitable combinations with especially low potential curvatures such as Ba+-LiNa.
@@
In these examples, the energy scale for s-wave collisions is comparable to typical ion-atom mixtures [30, 1]. There, the currently best experiments in rf-based hybrid traps are capable of mitigating micromotion-induced heating to a point where the collision energies reach $E_{col}/k_{B}\sim 10~{}\mu\text{K}$ which approaches the s-wave limit for the most favorable combinations. On the other hand, optical hybrid traps in principle allow for sympathetic cooling of any suitable ionic species to the temperature of the neutral ensemble. Assuming thermalization between a Ba+ ion and molecular quantum gases with now reachable temperatures of $\sim 10~{}\text{nK}$ [55, 57], this means that, for example, $E_{col}\approx E^{\star}_{J}/5$ for NaK and $E_{col}\approx E^{\star}_{J}/22$ for LiNa could be achieved, reaching into the quantum dominated regime. For this choice of the rotational state, ion-molecule experiments with collision energies on the order of $E_{J}^{\star}$ would likely require the capability to optically trap the ions and the molecules at the same time. On the other hand, studies of cold and ultracold chemical reactions and their additional control obtained by manipulating rotational degrees of freedom, which are at the focus of Section IV, or other experiments involving $E_{col}/k_{B}\gtrsim\text{mK}$ could be carried out using rf-based hybrid traps. In the following Section, we discuss the methodological requirements and feasibility of such hybrid ion-molecule platforms with respect to the currently available approaches and techniques.
III Experimental toolbox for hybrid trapping of ions and neutral molecules
Figure 3: (a) Schematic of an ion-molecule hybrid trap combining a magnetic trap, optical dipole traps, and an integrated trap for ions for investigating and controlling the interactions between single ions and neutral molecular quantum gases. (b) An ion immersed into the neutral cloud introduces a source of static electric fields, aligning the permanent molecular electric dipoles, depending on the distance $R$. This provides access to a new system with a rich hierarchy of competing tunable interactions spanning 1/$R^{6}$ (induced dipole – dipole, attractive), 1/$R^{4}$ (charge – induced dipole, attractive or repulsive: depending on the quantum numbers $J,m_{J}$), 1/$R^{3}$ (dipole – dipole, attractive for side-by-side or repulsive for head-to-tail collisions), 1/$R^{2}$ (charge – dipole, attractive or repulsive: determined by the electric field and $J,m_{J}$). The transition from the 1/$R^{4}$ to the 1/$R^{2}$ potential occurs above the characteristic electric field $\mathcal{E}^{\star}$.
In parallel to these developments in the field of ultracold molecular gases, first experiments for studying ultracold ion-atom collisions have been realized [13, 14]. Their aim was to combine vastly different methodologies for preparing and controlling atoms and ions at ultralow temperatures in a common two-species apparatus. Such hybrid setups using radiofrequency (rf)-driven Paul traps to confine ions and (magneto-) optical traps for preparing and controlling ultracold atomic ensembles have proven to be a versatile platform and have been successfully demonstrated for many ion-atom combinations [13, 14, 73, 15, 74, 75, 17, 26, 76, 77, 78, 79, 19, 24, 18]. More recently, it was shown that this concept can be extended to combinations of several species of atomic ions and multi-species atomic magneto-optical traps which can be used to study ion-atom collisions at low temperatures and in principle support the association of bialkali molecules [79]. Among the many important developments in this field, we point out two specific examples that are most relevant in view of the proposed extension of the study of ion-neutral interactions to ensembles of molecular quantum gases, that is, highly optimized rf-based hybrid traps and all-optical ion-atom traps. In the former case, for favorable ion-atom combinations featuring low reduced masses and long-range induction coefficients, optimized Paul traps in conjunction with advanced schemes for the detection and compensation of stray electric fields have allowed to reduce the collision energy to a range where non-classical effects become visible [24, 19]. In the latter case, similarly optimized setups were also combined with bichromatic optical traps capable of simultaneously confining ions and atoms [23, 34]. In particular, using optical traps provides the advantage that micromotion-induced heating can be completely eliminated and that this method is applicable to generic combinations of species provided that an effective optical trapping potential can be engineered [80]. In principle, this is the case for any combination of ions and neutral atoms or molecules possessing a non-zero dipole moment.
Therefore, we expect that a similar approach to overlapping ions and neutral species, that is, either using rf-based or all-optical hybrid traps, can also be applied to neutral molecules. In addition to ensembles of bialkali molecules such as those given in Table 1, this should also be the case for experiments on the level of individual particles, e.g. single molecules in optical tweezers [81, 82, 83, 84], other classes of polar molecules such as directly laser-coolable CaH [85], CaF [84], SrF [86], BaF [87], and YO [88], or polyatomic molecules like, e.g., CaOH [83] and CaOCH3 [89]. For clarity, we will focus our discussion on the exemplary ensembles of ground-state NaK and Ba+ ions for the following reasons. Firstly, molecular quantum (-degenerate) gases with a high phase-space density have been recently achieved for NaK [90, 55], and secondly, both the ions and the molecules can be routinely confined and prepared in standard optical dipole traps [91, 92, 93, 94, 90, 55]. Furthermore, as we will discuss in the remainder of this Section, NaK molecules prepared in the rovibronic ${}^{1}\Sigma^{+}$ ground state have very similar properties to common alkali species regarding mass and dynamic dipole polarizability $\alpha_{d}(\omega)$, with $\omega$ denoting the frequency of the interacting optical field. This allows for detailed comparisons of the expected behavior of the proposed ion-molecule system with Ba+-Rb mixtures which have been extensively studied both in presence of micromotion-induced heating encountered in rf-hybrid [25, 27, 26] and in rf-free optical traps where micromotion is eliminated [23]. Lastly, near-infrared (NIR) far-off resonant optical traps operated at a wavelength of 1064 nm, have been successfully used both to trap NaK and to demonstrate a generic approach to studying ion-atom collisions, first realized for Ba+-Rb mixtures [23]. To this end, bichromatic optical potentials are created by overlapping dipole traps of different wavelengths.
For a direct comparison with previous experiments we use the parameters employed in [23], that is, optical dipole traps with wavelengths of 532 nm (VIS ODT) and 1064 nm (NIR ODT), optical powers $P_{VIS}=0.47~{}\textrm{W}$ and $P_{NIR}=0.13~{}\textrm{W}$ (adjusted for typical temperatures of NaK quantum gases), matched $1/e^{2}$ waist radii $w_{0}^{VIS}=w_{0}^{NIR}=3.8\times 10^{-6}~{}\textrm{m}$, a residual stray electric field of $\mathcal{E}_{str}=10~{}\textrm{mV}~{}\textrm{m}^{-1}$, and the axial secular frequency of the linear Paul trap of $\omega_{x}/(2\pi)=12~{}\textrm{kHz}$, respectively. We then calculate the bichromatic potential suitable for overlapping a barium ion laser-cooled close to the Doppler temperature $T_{D}^{Ba}=365~{}\mu\textrm{K}$ with an ensemble of ground-state NaK molecules prepared at $T^{NaK}=400~{}\textrm{nK}$. The NaK molecules experience a repulsive potential from the VIS ODT due to its blue detuning with respect to the energetically lowest dipole-allowed optical transition $X^{1}\Sigma^{+}\leftrightarrow A^{1}\Sigma^{+}$. To maintain the initial trap depth $U_{0}^{NaK}/(k_{B})\approx 4~{}\mu\textrm{K}$, the intensity of the NIR ODT has to be increased compared to its initial value in order to achieve the desired differential light shift. Despite the substantially more complex energy level structure of molecules in comparison with alkali atoms, the resulting potentials calculated in the far-off-resonant limit [95] and using the dynamic dipole polarizability of NaK [60], that are shown in Fig.4 are quantitatively very similar to those obtained for the previously demonstrated case of Ba+-Rb in terms of trap depths and ODT parameters. This suggests that the protocols for overlapping ions and atoms demonstrated in [23] as well as proposed methods for further improving the trapping performance and realizing more complex trapping configurations [93, 80, 94] are readily applicable to molecules. The expected compatibility to such an extent may seem surprising at first glance because, in general, neutral atoms and polar molecules may have vastly different polarizabilities. This is indeed the case for the static polarizability $\alpha_{0}$ which is a pivotal point of the proposed experiments and is discussed in detail in Section II. However, in the optical and near-infrared frequency domains where the detunings with respect to rotational transitions are several orders of magnitude larger than those for optical transitions, all $\alpha_{d}(\omega)$ for the bialkali molecules in the rovibronic ground state as listed in Table 1 vary by less than a factor of 2 and are of the same order of magnitude as those of common alkali atoms [60]. It is this robustness of the dynamic polarizability that facilitates the creation of molecules from individually prepared atomic species which are typically all confined in the same far-detuned optical dipole traps. At the same time, even within the same electronic and vibrational ground state manifold, i.e. $\left|\textit{X}^{1}\Sigma^{+},v=0\right\rangle$, $\alpha_{d}(\omega)$ exhibits a substantial dependence on the rotational quantum numbers $J$ and $m_{J}$ [96].
Figure 4: (a) Effective bichromatic radial (orthogonal to the ODT’s wavevector) potential (black solid line, enlarged by a factor of 100 for visibility) for NaK (in the rovibronic ground state $\left|\textit{X}^{1}\Sigma^{+},\textrm{v}=0,\textrm{J}=0\right\rangle$) exposed to a composite optical dipole trap superimposing ODTs at 1064 nm (NIR, red dashed-dotted line) and at 532 nm (VIS, green dashed line). (b) The corresponding potentials for Ba+ (in the electronic ground state $6^{2}S_{1/2}$). The asymmetry and negative curvature of the effective potential at positions $\gtrapprox 4~{}\mu\text{m}$ around the trap center are caused by a residual stray electric field and the axial electrostatic potential of the Paul trap (grey dotted line), respectively. The latter is kept active during the optical trapping phase causing radial defocussing. The Paul trap and ODT parameters are adopted from experiments with Ba+-Rb mixtures [23], accounting for the properties of NaK at a temperature of $T^{NaK}=400~{}\textrm{nK}$.
Exploiting the similarities between molecules and alkali atoms interacting with far-detuned optical fields, a prototypical experimental setup could be realized as shown in Fig.3. Its central features include (i) a Paul trap optimized for ion-neutral experiments, that is, optical access through high-numerical-aperture objectives beneficial for stray electric field compensation and focusing of the dipole trap beam(s) and the correspondingly designed ion-electrode spacing $(\sim 1~{}\textrm{mm})$, (ii) magnetic field coils required for the magneto-association of Feshbach molecules, as well as (iii) a bichromatic optical dipole trap. Similar Paul traps have been demonstrated and feature suppression of stray electric fields down to $\sim 10^{-3}~{}\textrm{V m}^{-1}$ [23, 19, 97, 94], very low anomalous heating rates [92], and typical trap depths on the order of $k_{B}\times 10^{4}~{}\textrm{K}$, which is sufficiently deep to recapture ionic products.
The considerations above suggest that the proposed hybrid platform is suitable for investigations of ultracold ion-molecule interactions, such as precise studies of controlled quantum chemistry and ultracold collision experiments with dense molecular gases which are discussed in the following Sections IV and V, respectively.
IV Reactive collisions and formation of molecular ions